Six forces lying in a plane and forming angles of 60degrees relative to one another are applied to the centre of a homogenous sphere with a mass m=6kg.These forces are radially outward and consecutively 1N,2N,3N,4N,5N and 6N.The acceleration of the sphere is -----------.
 
Explain the procedure to find out acceleration.
 
(Answer is 1m/s^2)

Grade:11

1 Answers

Khimraj
5 years ago
try to resolve force in opposite direction 
then resultant will have three force each of 3N at 0o, 60o, 120o
then resulatnt of it will have two force of 3N and 3\sqrt{3} at angle 90
then resultant will be \sqrt{(3\sqrt{3})^{2} + 3^{2}} = 6N
Then acceleration = 6/6 = 1m/s2
Hope it clears.................................